OpenStudy (anonymous):

Assign a formal charge to each atom of AsCl3

OpenStudy (anonymous):

As = 5+; Cl = 1- As = 5-; Cl = 7+ As = 0; Cl = 0 As = 3+; Cl = 1- As = 6+; Cl = 2-

OpenStudy (abb0t):

You know that chlorine carries a 1\(^-\) charge, but you have 3, so you have 3\(^-\)! Now, to make this neutral, that means that you must have a what charge for As to make the overall charge = 0?

OpenStudy (anonymous):

+3